| Operation Summary for January 6, 2012 - January 9, 2012 |
NuMI reported receiving 9.32E18 protons on target for the period from 1/2/12 to 1/9/12. |
|
Friday January 6 The day shift began with Operations monitoring the Main Injector (MI) and Linac, with Booster personnel conducting aperture scan studies, and with the MiniBooNE and NuMI experiments taking beam. During the day shift, Operations reported that two Booster RF stations (BRF4 and BRF11) tripped off a number of times due to the aperture scan study. At 8:52 AM, the NuMI beam permit tripped off for the first of two times in about 90 minutes due to the output of a Lambertson magnet power supply (LAM60) going low; it reset. Operations established beam to the Meson FTBF experiment T-958 at 10:33 AM. At 1:58 PM, Operations shut off the Booster beam for about 10 minutes to replace a TCLK fan-out box. During the evening shift, An Operations Specialist and an Instrumentation person worked on the QXR at 5:22 PM. Operations disabled beam to T-958 about 10 PM. A Linac person came in and worked on a Linac RF station’s (LRF4) sporadic gradient drops. Saturday January 7 The midnight shift began with Operations sending beam to NuMI and MiniBooNE. During the midnight shift, two Linac quadrupole power supplies (L:QPS108 and L:QPS104) tripped off; they reset. Operators performed a Booster injection tune at 4:04 AM. At 4:34 AM, BRF18 tripped off; Operators investigated, but could not get its modulator to stay on. Operators bypassed the station and notified a Booster RF person. At 7:36 AM, the NuMI beam permit tripped off due to bad beam positions; it reset. During the day shift, Operators loaded the Recycler with protons at 8:53 AM. A Booster RF person worked on BRF18 and brought it back online at 9:34 AM. At 11:07 AM, L:QPS108 tripped its breaker for the first of four times in three hours; it reset. During the evening shift, Operators reported at 4:34 PM that the NuMI beam permit tripped off; it reset. At 6:30 PM, three I- Source (the backup source) vacuum pumps tripped off; they all reset. At 6:37 PM, BRF18 tripped off on an overload and would not reset; Operators bypassed the station and called in a RF technician who reported that the driver screen had a short and would need an access to repair. At 9:45 PM, BRF17 tripped off twice in 11 minutes; it reset. Operators reported that LIQPS108 tripped off; it reset. Sunday January 8 The midnight shift began with Operations sending beam to NuMI and MiniBooNE. During the midnight shift, LRF4 tripped off at 1:01 AM; it reset. Operations reported at 1:47 AM that the H- Source’s output was dropping; Operators investigated and found that its pressure wasn’t stable. Operators made adjustments. At 3:41 AM, Operations reported that the Booster RF West Anode Power Supply was being held off by an LCW leak indication on BRF18. Operators tried to reset the interlock, but it didn’t help. Operators contacted a RF person who said that since the station was off it would be okay to disconnect the cable and reset the interlocks. At 3:59 AM, BRF17 tripped off for the first of many times over two and a half hours; Operators adjusted its anode program and the station stopped tripping off. During the day shift, a Linac person began an hour-long emittance study at 8:42 AM. At 10:07 AM, LRF5 tripped off for the first of two times in 20 minutes; it reset. Operations established beam to T-958 at 12:11 PM. Operators loaded fresh protons into the Recycler at 1:55 PM. During the evening shift, LRF3 tripped off at 3 PM; it reset, but a Linac person came in to make adjustments. At 7 PM, the H- Source’s output dropped; Operators made adjustments. LRF4 tripped off at 7:35 PM; it reset. Operations disabled the T-958 beam at 10:05 PM. At 10:07 PM, the H- Source’s output dropped; Operators again made adjustments. L:QPS108 tripped off at 11:34 PM; it reset. Monday January 9 The midnight shift began with Operations sending beam to NuMI and MiniBooNE. During the midnight shift, Operators tuned the 750 KeV line at 12:57 AM. At 3:11 AM, BRF17 tripped off for the first of four times in 14 minutes; it reset. At 3:46 AM, the Linac Buncher tripped off; it reset. At 4:12 AM, the H- Source’s output dropped again; Operators made adjustments. A Booster RF vacuum pump (B:IPRF22) tripped off at 7:35 AM and took down BRF11 & 12 along with it; everything reset. Operations Coordinator (OpsCo) The OpsCo will allow Booster a two-hour access this morning to repair BRF18. |
|
More Information Comments and Suggestions |
| last modified 01/09/2012 email Fermilab |
| Security, Privacy, Legal | Use of Cookies | |